# On manifolds with almost non-negative Ricci curvature and integrally-positive $k^{th}$-scalar curvature

### v1

@@ 5.1 Sufficient curvature conditions for (SNP)
Taking into account (44), to conclude, it is sufficient to show that
$N\leq c(K,n)\frac{\widetilde{\mathsf{Vol}}(\tilde{B}_{1}(\bar{p}))}{\mathsf{Vol}(B_{1}(p^{\prime}))}.$ (45)
For every $i\in\{1,\cdots,l\}$, consider the isometric copies of $A_{i}$ contained in $\pi^{-1}(A_{i})$ that intersect $\tilde{B}_{3}(\bar{p})$. By (42), these are more than $N$ for every $i$. In particular, $\tilde{B}_{3}(\bar{p})$ is covered by the union of at least $N$ isometric copies of $A_{i}$ as $i$ varies. Hence
$\widetilde{\mathsf{Vol}}(\tilde{B}_{3}(\tilde{p}))\geq\sum_{i=1}^{l}N\,\mathsf{Vol}(A_{i})=N\,\mathsf{Vol}(B_{1}(p^{\prime})).$
Condition (45) then follows using Bishop-Gromov’s inequality on $\tilde{M}$. ∎
@@
Let $v,\epsilon,\delta,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. Let $(M^{n},g,p)$ be a manifold satisfying one of the conditions (34), (35), (36). Let $(\tilde{M}^{n},\tilde{g},\tilde{p})$ be a Riemannian covering, with covering map $\pi:\tilde{M}\to M$ such that $\pi(\tilde{p})=p$. Then $(\tilde{M},\tilde{g}/5^{2},\tilde{p})$ satisfies the corresponding condition among (34), (35), (36), where the constants have been rescaled by factors depending only on the dimension $n$.
@@
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $R,D,\delta>0$ with $R\geq 10D$ satisfying the following. Let $(M^{n},g)$ be a manifold satisfying one of the conditions (38), (39), (40). Let $(\tilde{M},\tilde{g},\tilde{p})$ be a covering of $M$ and let $\pi:\tilde{M}\to M$ be the covering map. Then, both $(M,g)$ and $(\tilde{M},\tilde{g})$ satisfy the segment neighbourhood property (SNP) (see Definition 4.2) w.r.t. $R$ and $D$.
@@
The statement follows combining Proposition 5.4 and Corollary 5.2 (and taking into account Remark 4.4). ∎
5.2 Applications of (SNP) to large scale geometry of manifolds
@@
The previous result fails if we remove the lower bounds on the Ricci curvature in conditions (38) and (40), thanks to the gluing results in [GromovLawsonSurgery1, GromovLawsonSurgery2, WolfsonKricci]. Indeed, consider consider the gluing of four copies of $S^{2}\times[0,+\infty)$ to $S^{3}$ minus four disjoint small balls. By [WolfsonKricci], we can equip this space with a Riemannian metric $g$ with the following properties:
@@
    $g$ coincides with the metric on $S^{3}$ and with the metric on each of the branches $S^{2}\times[0,+\infty)$ outside of a small neighbourhoof of the surgery.
@@
At the same time, Theorem 5.6 fails on $M$. With a similar construction using [GromovLawsonSurgery1, GromovLawsonSurgery2] one shows that also the Ricci bound in (40) cannot be removed.
@@
Proof of 1. Let $C>0$ be the constant given by Theorem 5.6. If $X$ has more than one end, then it is contained in a $C$-neighbourhood of a line by Theorem 5.6 and cannot have a third end.
Proof of 2. Let $\tilde{M}$ the universal cover of $M$. By Proposition 5.4 and Theorem 5.8, the volume of balls in $\tilde{M}$ grows at most linearly, so that by Proposition 2.22, it holds $\mathrm{b}_{1}(M)\leq 1$.
Proof of 3. It is clear that if $M$ is compact and its universal cover $\tilde{M}$ is non-compact, then $\pi_{1}(M)$ is infinite. We next show the converse. Since $\pi_{1}(M)$ has infinitely many elements, the universal cover $\tilde{M}$ is not compact. Hence, it suffices to show that $M$ is compact.
@@
Consider a lift $\tilde{r}_{M}$ of $r_{M}$ to $\tilde{M}$ such that $\tilde{r}_{M}(0)=M(p,\gamma_{1})$. Clearly, $\tilde{r}_{M}$ is a ray in $\tilde{M}$. By (47) and Lemma 4.6 ($\tilde{M}$ trivially has the (SNP)) and our choice of $t_{0},t_{1},t_{2}$, there exists $t\geq 10C$ such that $\tilde{r}_{M}(t)$ is $2C$-close to either $r_{\tilde{M}}(t_{0})$ or $r_{\tilde{M}}(t_{2})$. Assume without loss of generality that
@@
Proof of 4. Let $\gamma\subset M$ be a loop such that $[\gamma]\in\pi_{1}(M)$ has infinite period. Let $\tilde{\gamma}$ be a lift of $\gamma$ to $\tilde{M}$, and let $\tilde{\gamma}^{\mathbb{N}}:[0,+\infty)\to\tilde{M}$ be a lift of the infinite curve $\sum_{i\in\mathbb{N}}\gamma$. Since $[\gamma]\in\pi_{1}(M)$ has infinite period, it holds

### v2

@@ 5.1 Sufficient curvature conditions for a manifold to be thin
By Bishop-Gromov’s inequality, it holds #$\#I\leq c(K,n)$. Setting $p^{\prime}\in\{p_{i}\}_{i\in I}$ such that
[…]
and combining with (43), the statement follows. ∎
@@
Let $v,\epsilon,\delta,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. Let $(M^{n},g,p)$ be a manifold satisfying one of the conditions (34), (35), (36). Let $(\tilde{M}^{n},\tilde{g},\tilde{p})$ be the universal covering, with covering map $\pi:\tilde{M}\to M$ such that $\pi(\tilde{p})=p$. Then $(\tilde{M},\tilde{g}/3^{2},\tilde{p})$ satisfies the corresponding condition among (34), (35), (36), where the constants have been rescaled by factors depending only on the dimension $n$.
@@
Let $v,\epsilon,L\in(0,+\infty)$, $n\in\mathbb{N}$, $s\in(0,1)$ be fixed. There exist $R,D,\delta>0$ with $R\geq 20D$ satisfying the following. Let $(M^{n},g)$ be a manifold satisfying one of the conditions (38), (39), (40). Let $(\tilde{M},\tilde{g},\tilde{p})$ be the universal covering of $M$ and let $\pi:\tilde{M}\to M$ be the covering map. Then, both $(M,g)$ and $(\tilde{M},\tilde{g})$ are $(R,D)$-thin (see Definition 4.2).
@@
The statement follows combining Proposition 5.4 and Corollary 5.2. ∎
5.2 Large scale geometry of thin manifolds
@@
The previous result fails if we remove the lower bounds on the Ricci curvature in conditions (38) and (40), thanks to the gluing results in [GromovLawsonSurgery1, GromovLawsonSurgery2, WolfsonKricci]. Indeed, consider the gluing of four copies of $S^{2}\times[0,+\infty)$ to $S^{3}$ minus four disjoint small balls. By [WolfsonKricci], we can equip this space with a Riemannian metric $g$ with the following properties:
@@
    $g$ coincides with the metric on $S^{3}$ and with the metric on each of the branches $S^{2}\times[0,+\infty)$ outside of a small neighbourhood of the surgery.
@@
At the same time, Theorem 5.6 fails on $M$. With a similar construction using [GromovLawsonSurgery1, GromovLawsonSurgery2], one shows that also the Ricci bound in (40) cannot be removed.
@@
Proof of 1. The claim follows by Corollary 5.2 and Corollary 4.18.
Proof of 2. Let $C>0$ be the constant given by Theorem 5.6. If $X$ has more than one end, then it is contained in a $C$-neighbourhood of a line by Theorem 5.6 and cannot have a third end.
Proof of 3. Let $\tilde{M}$ the universal cover of $M$. By Proposition 5.4 and Theorem 5.8, the volume of balls in $\tilde{M}$ grows at most linearly, so that by Proposition 2.22, it holds $\mathrm{b}_{1}(M)\leq 1$.
Proof of 4. It is clear that if $M$ is compact and its universal cover $\tilde{M}$ is non-compact, then $\pi_{1}(M)$ is infinite. We next show the converse. Since $\pi_{1}(M)$ has infinitely many elements, the universal cover $\tilde{M}$ is not compact. Hence, it suffices to show that $M$ is compact.
@@
Consider a lift $\tilde{r}_{M}$ of $r_{M}$ to $\tilde{M}$ such that $\tilde{r}_{M}(0)=M(p,\gamma_{1})$. Clearly, $\tilde{r}_{M}$ is a ray in $\tilde{M}$. By (45), Lemma 4.5, and our choice of $t_{0},t_{1},t_{2}$, there exists $t\geq 10C$ such that $\tilde{r}_{M}(t)$ is $2C$-close to either $r_{\tilde{M}}(t_{0})$ or $r_{\tilde{M}}(t_{2})$. Assume without loss of generality that
@@
Proof of 5. Let $\gamma\subset M$ be a loop such that $[\gamma]\in\pi_{1}(M)$ has infinite period. Let $\tilde{\gamma}$ be a lift of $\gamma$ to $\tilde{M}$, and let $\tilde{\gamma}^{\mathbb{N}}:[0,+\infty)\to\tilde{M}$ be a lift of the infinite curve $\sum_{i\in\mathbb{N}}\gamma$. Since $[\gamma]\in\pi_{1}(M)$ has infinite period, it holds
@@
Conflict of interest statement. The authors have no relevant financial or non-financial interests to disclose. The authors have no conflict of interest to declare that are relevant to the content of this article.
Data Availability. Data sharing is not applicable to this article as no new data were created or analyzed in this study.